OpenStudy (anonymous):

I will give a medal, PLEASE HELP ME! I understand part one and part two, but I am having trouble with the third part asking if the graphs are the same and to explain my answer: Show that the solution of dy/dx = - (1+y^2)/(1+x^2) when y(0) = -1 is tan inverse (x) + tan inverse (y) = - pi/4 AND Show that tan inverse (x) + tan Inverse (y) = -pi/4 can be written as : y = (x+1)/(x-1) So my problem is this third part - are the graphs of tan inverse (x) +tan inverse (y) = -pi/4 AND y=(x+1)/(x-1) the same?? EXPLAIN YOUR ANSWER....PLEASE HELP! Thanks!

OpenStudy (phi):

I wrote \[ \tan^{-1} x + \tan^{-1} y = -\frac{\pi}{4} \\ \tan^{-1} y = - \tan^{-1} x - \frac{\pi}{4} \\ y = \tan \left( - \tan^{-1} x - \frac{\pi}{4} \right) \] and plotted both curves.

OpenStudy (phi):

They look the same

OpenStudy (anonymous):

Do you think that is a good enough explanation, I mean I would think so, my teacher grades so stinkin hard...Everyone is failing :(

OpenStudy (phi):

I hope so. Please post here how it goes... I am curious if there is something more they want.

OpenStudy (anonymous):

WHy does my graphing calculator show the graphs are different? Just curious if I am doing something wrong :)

OpenStudy (phi):

what is the difference?

OpenStudy (anonymous):

It shows graphs in the 2nd and 4th quadrant as well

OpenStudy (anonymous):

open study keeps kicking me off, so I am trying to stay with you, fyi...sorry if Im not here when ever you get back to me. THANKS!!

OpenStudy (anonymous):

I figured it out, thanks so much for the help!